OpenStudy (anonymous):

Calc question: An isotope of neptunium (Np-240) has a half-life of 65 minutes. If the decay of Np-240 is modeled by the differential equation dy/dt= -ky, where t is measured in minutes, what is the decay constant k? (My teacher didn't want to spend class time on this stuff because she says it is Algebra 2)

OpenStudy (anonymous):

Any help would be much appreciated

OpenStudy (jadzia):

\(\frac{dy}{dx} = -kdt\\ ln|y|=-kt+c\\ y=e^ce^{-kt} \\ y=Qe^{-kt}, Q=e^c\\ y=Qe^{-kt}\\ \frac{1}{2}Q=Qe^{-k(65)}\) then solve for k you'll get approx. 0.0106.

OpenStudy (anonymous):

Whoa, its dy/dt at the beginning. not dy/dx

OpenStudy (anonymous):

Where did all of the e's and Q's come from

OpenStudy (anonymous):

This made sense to me, but I didn't know where the 2 came from or why the k became positive

OpenStudy (jadzia):

Q is the constant for an exponential function.. i just used it as a variable. But you don't need it because it will cancel out at the end. e and q came from the general exponential function equation since you are dealing with half lives. not sure how to type in latex so, if you understand the process, you'll know the missing symbols.. first line should be: \[\int\limits \frac{ dy }{ y }=\int\limits -kdt\] i used the eqn button (not sure if it will show properly)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

So why does the k become positive

OpenStudy (anonymous):

and where did the 2 in the ln come from?

OpenStudy (jadzia):

1/2 not 2,, as I mentioned before, you're calculating 1/2 life. I assume you're familiar with the general formula for half life right? that formula is based from exponential function, which I used and derived it to calculate the half life last part: \(\frac{-1}{65} ln \frac{1}{2} = k\) evaluate and you'll get k

OpenStudy (anonymous):

Yeah. I'm a bit familiar with it

OpenStudy (anonymous):

So they just divided both sides by -1?

OpenStudy (anonymous):

And used the rules of logs?

OpenStudy (anonymous):

Ah I see now

OpenStudy (anonymous):

That's really sneaky :3

OpenStudy (jadzia):

yep

OpenStudy (anonymous):

Ty for the help

OpenStudy (jadzia):

no problem

OpenStudy (perl):

@Jadzia question, i have often wondered about raising a lower to the absolute value e^ln|y| = y , but shouldnt it be |y| ?

OpenStudy (jadzia):

yes it should be equal to |y|